# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ II-D Coding for Exchangeable Sources
If the source is also an i.i.d. mixture, then matching the codebook and the source distributions allows us to bound the overhead to $O(\log(k))$. However, for general exchangeable source distributions that are not i.i.d. mixtures, one can no longer simply set the codeword distribution to be the source distribution. In this section, we develop a novel code construction to allow a generalization of the result of Corollary 1 from i.i.d. sources to sources with arbitrary exchangeable distribution $p({\mathbf{x}})$.
One may be tempted to use the marginal distribution $p(x_{1})$ of the joint distribution of the source $p(x_{1},\cdots,x_{k})$ to generate i.i.d. codewords in the codebook. But this is not necessarily a good strategy. Consider an example where $(X_{1},\cdots,X_{k})$ takes on only two values, either $(0,\cdots,0)$ or $(1,\cdots,1)$ with probability 0.5 each. Using the marginal distribution would have resulted a Ber$(\frac{1}{2})$ codebook for which finding a matching index would have required searching over $O(2^{k})$ codewords, resulting in a common message rate of $O(k)$, whereas the optimal codebook should just have two codewords, i.e., $\{(0,\cdots,0),(1,\cdots,1)\}$ with a common message rate of only one bit.
The preceding example motivates us to define the following “urn” codebook for arbitrary exchangeable source distributions. Subsequently, we also defined “extended-urn” codebook for extendable source distributions.
@@ II-D1 Urn Codebook
The challenge lies in constructing length-$n$ codewords whose arbitrary subsequences of length $k$ all “look like” $\mathbf{x}$. Recall from Theorem 1 that the achievable rate for communicating a general exchangeable source $\mathbf{X}=(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ is $H(\mathbf{X})$ plus an overhead of $D(p\|q)$, where $q$ is an i.i.d. mixture distribution used for codebook construction.
The question is then: For an arbitrary exchangeable source distribution $p({\mathbf{x}})$, how should we design an i.i.d. mixture $q({\mathbf{x}})$ such that $D(p\|q)$ is small? We answer this question by introducing and analyzing the following novel codebook construction.
@@
Given an exchangeable source $\mathbf{X}$ with distribution $p({\mathbf{x}})$, an urn codebook $\mathbf{m}_{\text{URN}}=\left(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots,\right)$ is a codebook consisting of codewords $\mathbf{c}^{(j)}\in\mathcal{X}^{n}$ generated in the following fashion:
@@
    Sample a realization of $\mathbf{x}=(x_{1},\cdots,x_{k})$ using $p(\mathbf{x})$.
@@
    $\hat{p}_{\mathbf{x}}(c)=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{\{c\}}x_{i}.$ (23)
It is easy to see that the urn codebook is constructed from a mixture of i.i.d. distributions. In particular, the mixture weights are given by the probability that the sampled realization is of a particular type. Further, we argue that this i.i.d mixture is close to the source distribution. To see this, note that this codebook generation process can be alternatively viewed as repeatedly sampling with replacement from an urn containing the elements of $\mathbf{x}$. On the other hand, if the $k$ entries of a codeword are sampled without replacement from the entries of $\mathbf{x}$, then the $k$ entries would look as if they are distributed as $p(\mathbf{x})$. Thus, the difference between the source distribution and the i.i.d. mixture is precisely the difference between sampling with and without replacement. This intuition is captured in the following theorem, which quantifies this difference by bounding the KL divergence between the source distribution and the i.i.d. mixture induced by the code construction.
@@
$D(p\|q)\leq\min\{k\log(e),|\mathcal{X}|\log(k+1)\}.$ (24)
@@
See Appendix B
The above theorem gives upper bounds on the cost of using a codebook generated from an i.i.d. mixture distribution to represent an exchangeable source. The above bound immediately leads to the following achievability result for coded downlink massive random access.
@@
Consider a massive access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a discrete set $\mathcal{X}^{k}$ and be distributed according to an exchangeable distribution $p(\mathbf{x})$. Then the minimum common message rate $R^{*}$ for communicating each $X_{i}$ to its corresponding user $i$ is bounded above as
$R^{*}\leq H(\mathbf{X})+k\log(e)+\log(H(\mathbf{X})+k\log(e)+1)+1$ (25)
@@
$\begin{split}R^{*}&\leq H(\mathbf{X})+|\mathcal{X}|\log(k+1)\\ &\qquad\quad+\log(H(\mathbf{X})+|\mathcal{X}|\log(k+1)+1)+1.\end{split}$ (26)
@@
The implication is that for any exchangeable source, the overhead beyond $H({\mathbf{X}})$ is at most $O(k)$, or in case of finite alphabet sources $O(\log(k))$.
@@
It is interesting to note that the $O(k)$ scaling in overhead can already be obtained using a scheduling code of [6] to order the randomly activated $k$ users using at most $(k+1)\log(e)$ bits, followed by entropy coding of the sources listed in that order, as mentioned earlier. This approach of scheduling followed by source coding is different from the strategy of using the urn codebook, but the two achieve about the same $k\log e$ overhead for general exchangeable sources. The urn codebook can be thought of as combining scheduling and source coding steps together. The urn codebook also achieves a smaller overhead in case of finite alphabet sources.
@@ II-D2 Extended Urn Codebook
Although i.i.d. and i.i.d. mixture sources are also exchangeable, they do not incur the $O(k)$ overhead. The reason is that i.i.d. and i.i.d. mixtures are infinitely extendable. In this section, we investigate coding for exchangeable sources that are finitely extendable, and show that a better rate can already be achieved in this case. Recall that a source $(X_{1},\cdots,X_{k})$ is $d$-extendable if it has the same distribution as the first $k$ elements of a longer exchangeable sequence $(X_{1},\cdots,X_{k},\cdots,X_{d})$. The idea is that for a $d$-extendable source, we can construct modified version of the urn codebook by sampling from realizations of $(X_{1},\cdots,X_{k},\cdots,X_{d})$.
@@
Given an exchangeable source $(X_{1},\cdots,X_{k})$ which is $d$-extendable, i.e., there is an extended version $(X_{1},\cdots,X_{k},\cdots,X_{d})$ with distribution $p(x_{1},\cdots,x_{d})$ that is exchangeable, an extended urn codebook $\mathbf{m}_{\text{EX-URN}}=\left(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots,\right)$ is a codebook consisting of codewords $\mathbf{c}^{(j)}\in\mathcal{X}^{n}$ generated in the following fashion:
@@
    $\hat{p}_{\mathbf{x}}(c)=\frac{1}{d}\sum_{i=1}^{k}\mathbbm{1}_{\{c\}}x_{i}.$ (27)
@@
Finite de Finetti theorems are a class of results that provide bounds on some measure of distance between an i.i.d. mixture distribution (which is how the codewords in the extended urn codebook are constructed) and an exchangeable and $d$-extendable distribution. We discuss finite de Finetti results in detail in Section III. The following is a finite de Finetti theorem under KL divergence. It allows a characterization of the minimum common message rate for exchangeable and $d$-extendable sources.
@@
Consider a massive access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a discrete set $\mathcal{X}^{k}$ and be distributed according to an exchangeable and $d$-extendable distribution $p(\mathbf{x})$ with $d>k$. The minimum common message rate $R^{*}$ for communicating each $X_{i}$ to the respective user $i$ is bounded above as
$\begin{split}R^{*}&\leq H(\mathbf{X})+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)\\ &\qquad\quad+\log\left(H(\mathbf{X})+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)+1\right)+1\end{split}$ (29)
@@
$\begin{split}R^{*}&\leq H(\mathbf{X})+(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)\\ &\hskip 0.85358pt+\log\left(H(\mathbf{X})+(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)+1\right)+1.\end{split}$ (30)
@@
Observe that the first bound in (28) essentially reduces to the $k\log e$ bound in (24) in the case of $d=k$. For general exchangeable sources, this bound is tight. Theorem 6 later in the paper shows an example that achieves this bound. The second terms in (24) and (28) are the respective upper bounds in the case of finite alphabet sources. These bounds are derived using different proof techniques; the latter does not reduce to the former in the case of $d=k$.
@@ II-E Applications
II-E1 I.I.D. Source
As already shown in Corollary 1, when the source has i.i.d. or i.i.d. mixture distribution, the minimum downlink common message rate is the entropy of the source, plus an overhead of at most $O(\log(k))$. This overhead can be reduced to $O(1)$ if the source distribution is i.i.d. uniform.
@@ II-E2 Scheduling
Consider the problem of communicating a particular schedule to the $k$ active users where each user is only interested in learning its own slot assignment. For the case of $b=k$, the urn codebook would generate codewords according to uniform distribution on $[k]$. Therefore, we can use (18) in Theorem 1 to show that the following rate is achievable
$\begin{split}R^{*}&\leq H(\mathbf{X})+D(p\|q)+3\\ &\leq\log(k!)+k\log(e)+3\\ &\approx k\log(k),\end{split}$ (31)
@@
When $b>k$, this is an example of an exchangeable source distribution $p(\mathbf{x})$ that is $b$-extendable, because we can think of the process of scheduling $k$ users into $b$ slots as the first $k$ steps in the scheduling of $b$ users into $b$ slots (which is an extended exchangeable distribution).
Using the extended urn codebook, Corollary 3 allows us to bound the minimum common message rate as
$\begin{split}R^{*}&\leq H(\mathbf{X})+D(p\|q)+3\\ &\leq\log\left({{b}\choose{k}}k!\right)+\log\left(\frac{b^{k}}{b^{\underline{k}}}\right)+3\\ &\approx k\log(b)\end{split}$ (32)
The achievable rates in (31) and (32) have an intuitive explanation, as scheduling a user into one of the $b$ slots requires $\log(b)$ bits; so the minimum common message rate for scheduling $k$ users should scale as $k\log(b)$. It is remarkable that this common message rate is achieved using a code that does not specify exactly which user is scheduled into each slot.
A converse for these rates, which is tight to within a constant at large $n$, for the general case of $b\geq k$, can be obtained as follows:
@@
where $n$ is the total number of users. The proof uses a volume bound technique and is presented in Appendix D. For any fixed $k$, the second term vanishes as $n\to\infty$. Therefore, for large $n$, the achievable rate (31) and (32) are optimal to within a constant.
In the context of scheduling for massive random access, a different problem is studied in [6], where the BS can choose any schedule for the active users, as long as the slot assignments are non-colliding. In this case, assuming $b=k$, the common message rate can be further reduced by a $\log(k!)$ factor, because any of the $k!$ non-colliding schedules are equally allowable. This results in a common message rate of at most $(k+1)\log e$ bits as established in [6] and in agreement with the scaling in (31), where $R^{*}-\log(k!)\approx k\log e$.
@@ II-E3 Categorization
Assuming that all category assignments satisfying the constraint $\sum_{\ell=1}^{c}k_{\ell}=k$ are equally likely, the source $(X_{1},\cdots,X_{k})$, which represents the category labels for the $k$ active users, is an exchangeable random sequence. This exchangeable random sequence can also be thought of as drawing $k$ balls from an urn without replacement, where the urn contains $k$ balls in total; the balls are labeled with $\ell\in[c]$; and there are $k_{\ell}$ balls with label $\ell$.
To communicate the category labels to the active users, we can use an urn codebook, where the entries of the codewords can be thought of as drawing $k$ balls from the same urn, but with replacement. Notice that the urn contains $k_{\ell}$ balls for each label $\ell$ with $k_{\ell}$ fixed, so the codebook distribution $q(\mathbf{x})=\prod_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)^{k_{\ell}}$ is a constant over the support of $p(\mathbf{x})$. So, we can apply (18) in Theorem 1 to show that a rate of
$R^{*}\leq H(\mathbf{X})+D(p\|q)+3$ (34)
is achievable. By Theorem 2, the overhead $D(p\|q)$ is at most $\min\{k\log(e),c\log(k+1)\}$.
The above achievable common message rate can be computed more directly by expanding out the entropy terms:
$\begin{split}H(\mathbf{X})+D(p\|q)&=\sum_{\mathbf{x}\in[c]^{k}}p(\mathbf{x})\log\left(\frac{1}{q(\mathbf{x})}\right)\\ &=\log\left(\frac{1}{\prod_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)^{k_{\ell}}}\right)\\ &=k\sum_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)\log\left(\frac{k_{\ell}}{k}\right)=kH(\rho),\end{split}$ (35)
where $\rho=\left(\frac{k_{1}}{k},\cdots,\frac{k_{c}}{k}\right)$ is the empirical distribution of the labels. Thus interestingly, a characterizaton of achievable common message rate is simply the entropy of the category sizes, i.e.,
$R^{*}\leq kH(\rho)+3.$ (36)
@@
where $n$ is the total number of users. The proof is in Appendix E. Again, for any fixed $k$, the second term vanishes as $n\to\infty$. Therefore, for large $n$, the achievable rate (36) is optimal to within a constant.

### v4

@@ II-D Coding for Exchangeable Sources
If the sources are also i.i.d. mixture, then matching the codebook and the distribution of the sources allows us to bound the overhead on the common message length to $O(\log(k))$. However, for general exchangeable sources which are not i.i.d. mixtures, one can no longer simply set the codeword distribution to be the distribution of the sources. In this section, we develop a novel code construction to allow a generalization of the result of Corollary 1 from i.i.d. sources to sources with arbitrary exchangeable distribution $p({\mathbf{x}})$.
One may be tempted to use the marginal distribution $p(x_{1})$ of the joint distribution of the sources $p(x_{1},\cdots,x_{k})$ to generate i.i.d. codewords in the codebook. But this is not necessarily a good strategy. Consider an example where $(X_{1},\cdots,X_{k})$ takes on only two values, either $(0,\cdots,0)$ or $(1,\cdots,1)$ with probability $0.5$ each. Using the marginal distribution would have resulted a Ber$(\frac{1}{2})$ codebook for which finding a matching index would have required searching over $O(2^{k})$ codewords, resulting in a common message length of $O(k)$, whereas the optimal codebook should just have two codewords, i.e., the all-zero and the all-one vectors over the $n$ users, which would yield a common message length of only one bit.
The preceding example motivates us to define the following urn codebook for arbitrary exchangeable sources. Subsequently, we define extended-urn codebook for exchangeable sources which are extendable.
@@ II-D1 Urn Codebook
The challenge lies in constructing length-$n$ codewords whose arbitrary subsequences of length $k$ all “look like” $\mathbf{x}$. Recall from Theorem 1 that the achievable common message length for communicating general exchangeable sources $\mathbf{X}=(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ is $H(\mathbf{X})$ plus an overhead of $D(p\|q)$, where $q$ is an i.i.d. mixture distribution used for codebook construction.
The question is then the following. For arbitrary exchangeable sources with distribution $p({\mathbf{x}})$, how should we design an i.i.d. mixture $q({\mathbf{x}})$ such that $D(p\|q)$ is small? We answer this question by introducing and analyzing the following novel codebook construction.
@@
Given exchangeable sources $\mathbf{X}$ with distribution $p({\mathbf{x}})$, an urn codebook $\mathbf{m}_{\text{URN}}=\left(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots\right)$ is a codebook consisting of codewords $\mathbf{c}^{(j)}\in\mathcal{X}^{n}$ generated in the following fashion:
@@
    Generate a realization of $\mathbf{x}=(x_{1},\cdots,x_{k})$ according to the distribution $p(\mathbf{x})$.
@@
    Generate each entry of $\mathbf{c}^{(j)}$ in an i.i.d. fashion according to $\hat{p}_{\mathbf{x}}$, where $\hat{p}_{\mathbf{x}}$ is the empirical distribution (or the type) of $\mathbf{x}$, i.e.,
    $\hat{p}_{\mathbf{x}}(c)=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}(x_{i}=c),\quad\forall c\in\mathcal{X}.$ (33)
It is easy to see that the urn codebook is constructed from a mixture of i.i.d. distributions. In particular, the mixture weights are given by the probability that the sampled realization is of a particular type. Further, we argue that this i.i.d mixture is close to the distribution of the sources. To see this, note that this codeword generation process can be alternatively viewed as repeatedly sampling with replacement from an urn containing the elements of $\mathbf{x}$. If instead, every subset of $k$ entries in a codeword can all be viewed as if they are generated by sampling without replacement from the entries of $\mathbf{x}$, then the entries of the codeword would be distributed as $p(\mathbf{x})$. Thus, the difference between the exchangeable source distribution and the i.i.d. mixture codebook distribution is precisely the difference between mixtures of sampling with and without replacement. This intuition is captured in the following theorem, which quantifies this difference by bounding the KL divergence between the distribution of the sources and the i.i.d. mixture induced by the code construction.
@@
$D(p\|q)\leq\min\left\{\log\left(\frac{k^{k}}{k!}\right),|\mathcal{X}|\log(k+1)\right\}.$ (34)
The first term in the min operation in the above expression is further bounded from above by $k\log(e)$.
@@
See Appendix B. ∎
The above theorem gives upper bounds on the cost of using a codebook generated from an i.i.d. mixture distribution to represent sources which are exchangeable. The above bound immediately leads to the following achievability result for coded downlink massive random access.
@@
Consider a massive access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a discrete set $\mathcal{X}^{k}$ and be distributed according to an exchangeable distribution $p(\mathbf{x})$. Then the minimum common message length $R^{*}$ for communicating each $X_{i}$ to its corresponding user $i$ is bounded from above as
$R^{*}<H(\mathbf{X})+k\log(e)+\log(H(\mathbf{X})+k\log(e)+1)+2$ (35)
@@
$R^{*}<H(\mathbf{X})+|\mathcal{X}|\log(k+1)\\ \qquad\quad+\log(H(\mathbf{X})+|\mathcal{X}|\log(k+1)+1)+2.$ (36)
@@
The implication is that for exchangeable sources, the overhead beyond $H({\mathbf{X}})$ is at most $O(k)$, or in the finite alphabet case, at most $O(\log(k))$.
@@
It is interesting to note that the $O(k)$ scaling in overhead can already be obtained using a scheduling code of [6] to order the randomly activated $k$ users using at most $(k+1)\log(e)$ bits, followed by entropy coding of the sources listed in that order, as mentioned earlier. This approach of scheduling followed by source coding is different from the strategy of using the urn codebook, but the two achieve about the same $k\log(e)$ overhead for general exchangeable sources. The urn codebook can be thought of as combining scheduling and source coding steps together. The urn codebook also achieves a smaller overhead in case of finite alphabet sources.
@@ II-D2 Extended Urn Codebook
Although i.i.d. and i.i.d. mixture sources are also exchangeable, they incur less than $O(k)$ overhead. The reason is that i.i.d. and i.i.d. mixtures are infinitely extendable. In this section, we investigate coding for exchangeable sources that are finitely extendable, and show that a better common message length can already be achieved in this case. Recall that sources $(X_{1},\cdots,X_{k})$ are $d$-extendable if they have the same distribution as the first $k$ elements of a longer exchangeable sequence $(X_{1},\cdots,X_{k},\cdots,X_{d})$. The idea is that if the sources are $d$-extendable, we can construct modified version of the urn codebook by sampling from realizations of $(X_{1},\cdots,X_{k},\cdots,X_{d})$.
@@
Given exchangeable sources $(X_{1},\cdots,X_{k})$ which are $d$-extendable, i.e., there is an extended exchangeable sequence $(X_{1},\cdots,X_{k},\cdots,X_{d})$ with distribution $p(x_{1},\cdots,x_{d})$, an extended urn codebook $\mathbf{m}_{\text{EX-URN}}=\left(\mathbf{c}^{(1)},\mathbf{c}^{(2)},\cdots\right)$ is a codebook consisting of codewords $\mathbf{c}^{(j)}\in\mathcal{X}^{n}$ generated in the following fashion:
@@
    $\hat{p}_{\mathbf{x}}(c)=\frac{1}{d}\sum_{i=1}^{d}\mathbbm{1}(x_{i}=c),\quad\forall c\in\mathcal{X}.$ (37)
@@
Finite de Finetti theorems are a class of results that provide bounds on some measure of distance between an i.i.d. mixture distribution (which is how the codewords in the extended urn codebook are constructed) and an exchangeable and $d$-extendable distribution. We discuss finite de Finetti results in detail in Section III. The following is a finite de Finetti theorem under KL divergence. It allows a characterization of the minimum common message length for exchangeable and $d$-extendable sources.
@@
Consider a massive access scenario with a total of $n$ users and a random subset of $k$ active users. Let sources $\mathbf{X}=(X_{1},\cdots,X_{k})$ take values in a finite or countable set $\mathcal{X}^{k}$ and be distributed according to an exchangeable and $d$-extendable distribution $p(\mathbf{x})$ with $d>k$. The minimum common message length $R^{*}$ for communicating each $X_{i}$ to the respective user $i$ is bounded from above as
$R^{*}<H(\mathbf{X})+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)\\ +\log\left(H(\mathbf{X})+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)+1\right)+2$ (39)
@@
$R^{*}<H(\mathbf{X})+(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)\\ \qquad\qquad+\log\left(H(\mathbf{X})+(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)+1\right)+2.$ (40)
@@
Observe that the first bound in (38) essentially reduces to the $k\log(e)$ bound in (34) in the case of $d=k$. For general exchangeable sources, this bound is tight. Theorem 6 later in the paper shows an example that achieves this bound. The second terms in (34) and (38) are the respective upper bounds in the case of finite alphabet sources. These bounds are derived using different proof techniques; the latter does not reduce to the former in the case of $d=k$.
Remark 7
The finite alphabet-size bound in (38) and (39), i.e.,
$(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)$ (41)
is due to the analysis in [10] on the KL divergence of the probability distributions resulting from sampling with and without replacement. It is possible to replace this term by
$(|\mathcal{X}|-1)\frac{k(k-1)}{2(d-1)(d-k+1)}$ (42)
as in a similar analysis in [9]. We explore this connection further in Section III-B.
@@ II-E Applications
II-E1 I.I.D. Sources
As already shown in Corollary 1, when the sources are i.i.d. or i.i.d. mixture distribution, the minimum downlink common message length is the entropy of the sources, plus an overhead of at most $O(\log(k))$. This overhead can be reduced to $O(1)$ if the sources are i.i.d. and uniformly distributed.
@@ II-E2 Scheduling
Consider the problem of communicating a particular schedule to the $k$ active users where each user is only interested in learning its own slot assignment. For the case of $b=k$, the urn codebook would generate codewords according to uniform distribution on $[k]$. Therefore, we can use (26) in Theorem 1 to show that the following common message length is achievable
$R^{*}$ $<$ $H(\mathbf{X})+D(p\|q)+5$ (43)
$\leq$ $\log(k!)+k\log(e)+5$ (44)
$\approx$ $k\log(k),$ (45)
@@
When $b>k$, this is an example of exchangeable sources with distribution $p(\mathbf{x})$ that is $b$-extendable, because we can think of the process of scheduling $k$ users into $b$ slots as the first $k$ steps in the scheduling of $b$ users into $b$ slots (which is an extended exchangeable distribution).
Using the extended urn codebook, Corollary 3 allows us to bound the minimum common message length as
$R^{*}$ $<$ $H(\mathbf{X})+D(p\|q)+5$ (46)
$\leq$ $\log\left({{b}\choose{k}}k!\right)+\log\left(\frac{b^{k}}{b^{\underline{k}}}\right)+5$ (47)
$\approx$ $k\log(b).$ (48)
The achievable common message lengths in (45) and (48) have an intuitive explanation, as scheduling a user into one of the $b$ slots requires $\log(b)$ bits; so the minimum common message length for scheduling $k$ users should scale as $k\log(b)$.
A converse for these common message lengths, which is tight to within a constant at large $n$, for the general case of $b\geq k$, can be obtained as follows:
@@
where $n$ is the total number of users. The proof uses a volume bound technique and is presented in Appendix D. For any fixed $k$, the second term vanishes as $n\to\infty$. Therefore, for large $n$, the achievable common message lengths (45) and (48) are optimal to within a constant.
In the context of scheduling for massive random access, a different problem is studied in [6], where the BS can choose any schedule for the active users, as long as the slot assignments are non-colliding. In this case, assuming $b=k$, the common message length can be further reduced by a $\log(k!)$ factor, because any of the $k!$ non-colliding schedules are equally allowable. This results in a common message length of at most $(k+1)\log e$ bits as established in [6] and in agreement with the scaling in (45), where $R^{*}-\log(k!)\approx k\log e$.
@@ II-E3 Categorization
Assuming that all category assignments satisfying the constraint $\sum_{\ell=1}^{c}k_{\ell}=k$ are equally likely, the sources $(X_{1},\cdots,X_{k})$, which represent the category labels for the $k$ active users, form an exchangeable random sequence. This exchangeable random sequence can also be thought of as drawing $k$ balls from an urn without replacement, where the urn contains $k$ balls in total; the balls are labeled with $\ell\in[c]$; and there are $k_{\ell}$ balls with label $\ell$.
To communicate the category labels to the active users, we can use an urn codebook, where the entries of the codewords can be thought of as drawing $k$ balls from the same urn, but with replacement. Notice that the urn contains $k_{\ell}$ balls for each label $\ell$ with $k_{\ell}$ fixed, so the codebook distribution $q(\mathbf{x})=\prod_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)^{k_{\ell}}$ is a constant over the support of $p(\mathbf{x})$. So, we can apply (26) in Theorem 1 to show that the following common message length is achievable:
$R^{*}<H(\mathbf{X})+D(p\|q)+5.$ (50)
By Theorem 2, the overhead $D(p\|q)$ is at most $\min\{k\log(e),c\log(k+1)\}$.
The above achievable common message length can be computed more directly by expanding out the entropy terms:
$H(\mathbf{X})+D(p\|q)$ $=$ $\sum_{\mathbf{x}\in[c]^{k}}p(\mathbf{x})\log\left(\frac{1}{q(\mathbf{x})}\right)$ (51)
$=$ $\log\left(\frac{1}{\prod_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)^{k_{\ell}}}\right)$ (52)
$=$ $k\sum_{\ell=1}^{c}\left(\frac{k_{\ell}}{k}\right)\log\left(\frac{k_{\ell}}{k}\right)$ (53)
$=$ $kH(\rho),$ (54)
where $\rho=\left(\frac{k_{1}}{k},\cdots,\frac{k_{c}}{k}\right)$ is the empirical distribution of the labels. Thus interestingly, a characterization of achievable common message length is simply the entropy of the category sizes, i.e.,
$R^{*}<kH(\rho)+5.$ (55)
@@
where $n$ is the total number of users. The proof is in Appendix E. Again, for any fixed $k$, the second term vanishes as $n\to\infty$. Therefore, for large $n$, the achievable common message length (55) is optimal to within a constant.